A plane flies 400 miles with the wind and 320 miles against the wind in the same length of time. If the speed of the wind is 20 mph, what is the speed of the plane in still air?
Answer by mananth(16946)

A plane flies 400 miles with the wind and 320 miles against the wind in the same length of time. If the speed of the wind is 20 mph, what is the speed of the plane in still air?
let speed in still air be x mph
..
distance = 400
speed with wind = x+20
time = 400/(x+20)
...
against wind
distance = 320
speed = x-20
time = 320/(x-20)
...
both times are equal
400/(x+20)= 320/(x-20)
400(x-20)=320(x+20)
400x-8000=320x+6400
400x-320x=8000+6400
80x=14400
x=180 mph
...
m.ananth@hotmail.ca